### v1

# Prediction Instability in Machine Learning Ensembles

Abstract
In machine learning ensembles predictions from multiple models are aggregated. Despite widespread use and strong performance of ensembles in applied problems little is known about the mathematical properties of aggregating models and associated consequences for safe, explainable use of such models. In this paper we prove a theorem that shows that any ensemble will exhibit at least one of the following forms of prediction instability. It will either ignore agreement among all underlying models, change its mind when none of the underlying models have done so, or be manipulable through inclusion or exclusion of options it would never actually predict. As a consequence, ensemble aggregation procedures will always need to balance the benefits of information use against the risk of these prediction instabilities. This analysis also sheds light on what specific forms of prediction instability to expect from particular ensemble algorithms; for example popular tree ensembles like random forest, or xgboost will violate basic, intuitive monotonicity and fairness properties.
footnotetext: 1Department of Electrical Engineering and Computer Science, Milwaukee School of Engineering, Milwaukee, WI, USA. Correspondence to: Jeremy Kedziora <kedziora@msoe.edu>.
Preliminary work, Copyright 2024 by the author(s).
1 Introduction
The promise of machine learning and artificial intelligence is to automate decision making in an efficient, rational form that makes optimal use of information and offers consistency of choice. A popular strategy to optimize machine learning-based decision making is to maximize predictive accuracy by aggregating the insights of multiple models together, a process called ensembling. Decades of research and experimentation have not revealed an obvious “best” method for aggregating models, an outcome mirroring no-free-lunch theorems. Moreover, despite widespread use and strong performance of ensembles in applied problems little is known about the mathematical properties of aggregating models and associated consequences for safe, explainable use of such models. An alternative to focusing solely on maximizing predictive accuracy would be to proceed axiomatically by setting down desirable properties in an aggregation procedure and then characterizing the set of aggregation rules that satisfies them.
To motivate the axiomatic approach, consider a tree ensemble used for classification trained via e.g. the random forest algorithm [Breiman, 1996, 1999, 2001] or the xgboost algorithm [Chen and Guestrin, 2016]–a collection of $m$ decision trees with tree $j$ represented as a pair $(T_{j},L_{j})$ where $T_{j}$ is a partition of the data and $L_{j}(x)$ is a function mapping from an observation in the feature space $x$ to an element of the partition. One method to use this model to make predictions or choices would be to aggregate scores from the $m$ trees and choose the class $y$ with the largest such score:
$\operatornamewithlimits{argmax}_{y}\left\{\frac{1}{m}\sum_{j=1}^{m}\left(\frac{1}{|L_{j}(x)|}\sum_{y^{\prime}\in L_{j}(x)}I(y=y^{\prime})\right)\right\}$
where $I(\cdot)$ is the indicator function. As an example, suppose that we fit such an ensemble on a data set with three classes and then, for a pair of points in the feature space $x_{1}$ and $x_{2}$, conduct inference to obtain the following:
Scores for $x_{1}$:
Class tree 1 tree 2 tree 3 Agg. Score
1 0.40 0.40 0.40 0.4000
2 0.34 0.35 0.5 0.3967
3 0.26 0.25 0.1 0.2033
Scores for $x_{2}$:
Class tree 1 tree 2 tree 3 Agg. Score
1 0.40 0.40 0.40 0.4000
2 0.36 0.35 0.5 0.4033
3 0.24 0.25 0.1 0.1967
From this we would predict or choose class $1$ for $x_{1}$ and class 2 for $x_{2}$. But notice that these two points in feature space are similar in the sense that trees 2 and 3 have identical scores, and tree 1 yields that class 1 is more likely than class 2, which is more likely than class 3 in both cases. Despite this, after aggregation the ensemble prediction or choice is not the same for these two examples. That is, the ensemble “changed its mind” about the relative likelihood of the classes even though no individual model did.
In this note we argue that this sort of prediction or decision-making inconsistency is an intrinsic part of ensembling. Specifically, we prove a theorem that shows that any ensemble will exhibit at least one of the following forms of prediction instability. It will either ignore agreement among all underlying models, change its mind when none of the underlying models have done so, or be manipulable through inclusion or exclusion of options it would never actually predict. As a consequence, ensemble aggregation procedures will always need to balance the benefits of information use against the risk of choice instability. We formalize this tradeoff in what we will call an ensemble choice discontinuity theorem.
The remainder of this section lays out other work that we draw upon, the second section introduces our framework and axioms, the third specifies our main results, and the fourth discusses implications of these results and concludes.
Related Work
The argument for ensembling many models together can be traced back to the jury theorems inspired by de Condorcet [1785]. These methods are heavily used in classical supervised learning [e.g. Breiman, 2001], but also appear in modern deep learning [Ganaie et al., 2022], in the token generation of large language models [e.g. Shazeer et al., 2017], and in reinforcement learning [Song et al., 2023].
Modern use of ensembling techniques tends to involve choosing a base algorithm to construct learners, sampling data/features to train that base algorithm via bagging [Breiman, 1996, 1999, 2001] or boosting [Freund and Schapire, 1997] numerous times, and finally aggregating the outputs of individual models, e.g via hard or soft voting, or some learned procedure, e.g. stacking [Wolpert, 1992]. [^1] See Int [2000], Maclin and Opitz [2011], Polikar [2006], or Rokach [2010] for reviews. Despite the availability of numerous options for aggregation little scholarship has focused on studying the qualities of different aggregation procedures. As others [e.g. Wolpert and Macready, 1997, Werbin-Ofir et al., 2019] have noted, the optimal choice of aggregator is likely to be problem-specific. Given this, we study aggregators from an axiomatic perspective. This allows us to identify combinations of desirable properties, seek the aggregators that satisfy them, and thereby characterize the consistency and stability of the predictions or decision-making derived from ensembles.
Social choice theory has long analyzed aggregative mechanisms axiomatically in the context of economics and political science. One of the earliest social choice results, Arrow’s impossibility theorem argued that no method of aggregating the ordinal preferences of a set of individuals into a single group-level ordinal preference could satisfy a set of desirable properties chosen to maximize consistency in decision-making and democratic principles [see Arrow, 1951].
We take advantage of three specific extensions to Arrow’s original result to characterize ensemble aggregation. [^2] For a review see Patty and Penn [2019]. First, Arrow analyzed an orders-to-order mapping and machine learning models typically output scores, implying a scores-to-score mapping. Harsanyi [1955], Samuelson [1967], Sen [1970a], Schwartz [1970], Fishburn [1972], Kalai and Schmeidler [1977] and others extended Arrow’s result to cardinal preferences and we utilize their insights here. Second, in operational contexts machine learning is often used to make decisions, implying a focus on the choice derived from the scores-to-score mapping rather than the output score itself. This mirrors efforts to analyze the compatibility of Arrow’s axiomatic approach with so-called social choice functions that identify a subset of the alternatives rather than a preference ordering over them [e.g. Sen, 1969, 1970b, 1971] and that inform our own formalization. Finally, efforts have been made to show that Arrow’s axioms are special cases of other properties, leading to more general impossibility theorems. We will make use of Eliaz [2004] in this regard.
2 Framework
We begin by formalizing the notion of an ensemble. Consider a variable that an analyst wishes to predict with possible values $y$ given by a set $Y$ so that $y\in Y$. We will refer to this variable as the response and its values as labels. If it is a regression problem then $Y\subseteq\mathbb{R}$. If the goal is to model a count then $Y$ is countably infinite. If it is a classification problem then $Y$ is finite.
With this response variable we will associate a set of predictor variables referred to as features and assumed to be an element $x$ of some feature data space $X$. Given a data set of examples training a model $j$ is equivalent to choosing a scoring function $s_{j}:X\to S_{j}(Y)\subseteq\mathbb{R}^{|Y|}$ where $s_{j}(x)=\{s_{j}(x,y)\}_{y\in Y}$. Beyond that, let $\mathbf{R}$ denote the set of all complete and transitive binary relations on $Y$; any of these binary relations will induce an ordering of the label values in $Y$. Given a particular ordering $R\in\mathbf{R}$ we can collect all the different scores consistent with this ordering as a set:
$F(Y,R)=\left\{z\in\mathbb{R}^{|Y|}\left|\mbox{for all }y,y^{\prime}\in Y\atop{y\hskip 2.84526ptR\hskip 2.84526pty^{\prime}\mbox{ iff }z_{d(y)}\geq z_{d(y^{\prime})}}\right.\right\}$
where $d(y)$ is the dimension associated with the score for $y$. We assume throughout that for all $R\in\mathbf{R}$ we have $S_{j}(Y)\bigcap F(Y,R)\neq\emptyset$ so that the model is not precluded from encoding any ordering of the label values. [^3] Note that this is not the same as assuming surjectivity of $s_{j}(\cdot)$. We will have more to say about how this relates to our results in the discussion section.
This model scoring function maps the feature inputs to scores that can be used to predict values of $y$ accurately, according to some well-defined notion of accuracy. In analogue with our motivating tree ensemble example, a model scoring function could be the output of a single decision tree.
Model Ensembles
The intuition of ensembling is that it should be possible to leverage the complementary qualities and diversity of different models to create a more robust and accurate representation of $y$. Thus, we suppose that we have a set of $m\geq 2$ models, each with its own scoring function learned from the data indexed by $j$, and each mapping into $S_{j}(Y)\subseteq\mathbb{R}^{|Y|}$ so that:
$s(x)=$$\left\{\left\{s_{j}(x,y)\right\}_{y\in Y}\right\}_{j=1,\ldots,m}$
$\hskip 56.9055pt\in\prod_{j}S_{j}(Y)\equiv S(Y).$
In general, we will suppress dependence on $x$ unless we need to distinguish between different input features.
With a set of model scoring functions in hand we will model the ensemble prediction as a choice from the set of label values, $Y$. Accordingly, define an ensemble choice aggregator as a set-valued function: $C:Y\times S(Y)\rightarrow\mathcal{P}(Y)\backslash\emptyset$, where $\mathcal{P}(Y)$ is the power set of $Y$ [^4] We exclude $\emptyset$ from the range of $C(\cdot,s)$ to focus throughout on ensemble choice aggregators that actually make choices. Later, we will use this to rule out a form of intransitivity in Lemma 3.2. The job of this function is to use the model scores to make a choice of one or more options from $Y$. In general, we will assume for any $Y^{*}\subseteq Y$ and any $s$ that $y\in C(Y^{*},s)$ if there does not exist $y^{\prime}\in Y^{*}$ such that $y^{\prime}=C(\{y,y^{\prime}\},s)$.
Aggregator Properties
Our next task is to specify properties that seem like a priori desirable characteristics in an ensemble. Our goal will then be to assess the extent to which it is possible to have an ensemble score aggregator that satisfies them. We will focus on two ways to assess choice consistency:
  1. 1.
    How do the decisions made by an ensemble respond to post-training changes to the set of labels the response could take on?
  2. 2.
    How the decisions made by an ensemble respond when we impose requirements on how the score aggregator uses information from the underlying models?
Change to the set of possible labels comes in two forms: deletions and insertions. For example, an online retailer using a recommender system could choose to permanently drop products from its catalogue; if it is expensive to retrain models then deleting a label option would be a temporary method to ensure that recommendations remain in sync with available products. We propose the following definition–adapted to our context from Arrow [1959] and Sen [1969]–to formalize ensemble decision consistency given these types of changes:
Definition 2.1.
Say that $C(\cdot,s)$ is insertion/deletion consistent if for all $s\in S(Y)$ and any $Y_{1}^{*},Y_{2}^{*}\subseteq Y$ such that $Y_{1}^{*}\subseteq Y_{2}^{*}$ we have that $Y_{1}^{*}\bigcap C(Y_{2}^{*},s)\neq\emptyset$ implies that $C(Y_{1}^{*},s)=Y_{1}^{*}\bigcap C(Y_{2}^{*},s)$.
Insertion/deletion consistency requires that if the ensemble chooses or predicts some label options when all possible labels for the response are available and then the set of labels is reduced but still contains some previously chosen labels then 1) no label passed over before now becomes chosen and 2) no previously chosen label is now passed over. In other words, the ensemble cannot be compelled to change what it does choose via requiring that it ignore label options it never chose in the first place; adding new label options cannot break ties. [^5] Sen [1969] gives the following examples of this property for intuition on page 384 of his original paper: “…if the world champion in some game is a Pakistani, then he must also be the champion in Pakistan…if some Pakistani is a world champion, then all champions of Pakistan must be champions of the world.”
We turn next to consider desirable requirements on how an ensemble uses information, adapted from Arrow [1951], Sen [1970a], and Eliaz [2004].
Definition 2.2.
Say that $C(\cdot,s)$ is a nondegenerate ensemble choice aggregator if there does not exist $j$ such that for all $y,y^{\prime}\in Y$ and all $s\in S(Y)$ if $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ then $y=C(\{y,y^{\prime}\},s)$.
Definition 2.3.
Say that $C(\cdot,s)$ satisfies ensemble unanimity if for all $y,y^{\prime}\in Y$ and all $s\in S(Y)$ we have that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ for all $j$ implies $y=C(\{y,y^{\prime}\},s)$.
Definition 2.4.
Say that $C(\cdot,s)$ respects model choice reversal if for all $y,y^{\prime}\in Y^{*}\subseteq Y$ and all $s,s^{\prime}\in S(Y)$ we have that $y\in C(\{y,y^{\prime}\},s)$, $y^{\prime}\notin C(\{y,y^{\prime}\},s)$, and $y^{\prime}\in C(\{y,y^{\prime}\},s^{\prime})$ implies that there exists $j$ such that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ and $s^{\prime}_{j}(\cdot,y)<s^{\prime}_{j}(\cdot,y^{\prime})$.
First, if the intuition of ensembling–that many models improve on decision-making over a single model–is born out then the ensemble should use the information of multiple models to produce its decisions and the aggregator should not be an isomorphism of any single model. Definition 2.2 is a minimal version of such a requirement in that it requires that a single model does not completely determine the final choice. Second, it seems natural to assume that if every model in the ensemble scores $y$ higher than $y^{\prime}$ then the choice of the ensemble should reflect this; we formalize this in Definition 2.3. Third, how should the propensity of the ensemble to change its decision between labels depend on the propensity of the underlying models to do so? In Definition 2.4 we stipulate that the ensemble cannot change its mind for $y$ and $y^{\prime}$ unless at least one model in the ensemble changed its mind for $y$ and $y^{\prime}$. [^6] Model choice reversal is a generalization of the independence of irrelevant alternatives property used by Arrow [1951] and many others. See Eliaz [2004] Propositions 1 and 2 for analysis.
3 An Ensemble Choice Discontinuity Theorem
In this section we will establish our main result. We will argue that there is a fundamental tradeoff between information use and stability of ensemble predictions and choices in the sense that no aggregation procedure satisfies defintions 2.1, 2.2, 2.3, and 2.4, a result that we refer to in shorthand as ensemble choice discontinuity (ECD). Our focus here will be on building intuition for why our ECD theorem is true via sketching the arguments for it. The following section will expand on implications with examples and discussion. Full proofs for all results are in the appendix.
Central to the proof of our ECD theorem is the notion of transitivity:
Definition 3.1.
Say that $C(\cdot,s)$ satisfies transitivity if for all $y,y^{\prime},y^{\prime\prime}\in Y$ and all $s\in S(Y)$ we have that $y\in C(\{y,y^{\prime}\},s)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$ implies $y\in C(\{y,y^{\prime\prime}\},s)$.
In the first step of our argument we show that any ensemble choice aggregator that satisfies insertion/deletion consistency will be transitive, formalized in this Lemma:
Lemma 3.2.
$C(\cdot,s)$ satisfies insertion/deletion consistency if and only if it is transitive.
So, given an ensemble choice aggregator that is insertion/deletion consistent, and so transitive, the next step is to argue that transitivity implies that at least one information use requirement must be violated. The strategy to show this follows insights from Geanakoplos [2005] and Fey [2014]. The first step is to use ensemble unanimity to identify a set of scores in which there is a critical model – a candidate for a model in the ensemble whose scores might completely determine ensemble choices. The second part of the argument is to apply ensemble unanimity, transitivity, and model choice reversal to a series of scores derived from making “small” changes to the original set of scores used to identify the critical model. In particular, where model choice reversal holds, if no model reverses its scores for one label versus another then the ensemble cannot change its choices between the two labels. Formally:
Lemma 3.3.
Suppose that $C(\cdot,s)$ respects model choice reversal. Then $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ iff $s_{j}^{\prime}(\cdot,y)>s_{j}^{\prime}(\cdot,y^{\prime})$ for all $j$ implies that $C(\{y,y^{\prime}\},s)=C(\{y,y^{\prime}\},s^{\prime})$.
Lemma 3.3 allows us to show that each of the series of model scores must yield the same choice as a set of scores in which the critical model determines the choice; we conclude that $C(\cdot,s)$ must not be a nondegenerate ensemble choice aggregator, leading to our main result, which we now state:
Theorem (Ensemble Choice Discontinuity).
Suppose that $|Y|\geq 3$. If $C(\cdot,s)$ is a nondegenerate ensemble choice aggregator then it must violate at least one of insertion/deletion consistency, ensemble unanimity, or model choice reversal.
4 Implications
What does this theoretical result mean for how we should expect machine learning ensembles to behave in practical settings and what lessons do we take away? In general, the ECD theorem tells us that if we would like to have a machine learning ensemble use the insights of multiple models in making predictions and choices (definition 2.2) then we must accept the risk that the ensemble makes those predictions and choices in a counterintuitive or unstable fashion at least some of the time, ignoring model agreement (definition 2.3), changing its mind in a way that the underlying models cannot justify (definition 2.4) or in response to the presence of label options it would not predict or choose (definition 2.1).
Violations of insertion/deletion consistency are a form of discontinuity in the scoring space in the sense that changes to the unchosen portion of the set of label values can induce changes in the ensemble prediction or choice even if the individual model scores are held constant. Violations of model choice reversal are a form of discontinuity in the underlying information in the sense that two scores can both be consistent with the same ordering of the set of label values for all models and yet result in different ensemble predictions or choices.
We expand on this with some examples. To make this discussion concrete, a natural way to use an ensemble to make decisions about labels would be to aggregate the scores of individual models and then require that the ensemble choose the label that has the highest aggregated score. This method of making choices is clearly insertion/deletion consistent and transitive under any method of score aggregation.
In a voting ensemble individual model scores are combined to make a final prediction or decision through either hard or soft voting. In hard voting the voting ensemble predicts the class label that receives the plurality of votes from the constituent models in which case the ensemble choice aggregator will be:
$C(Y^{*},s)$
$\hskip 8.53581pt=\operatornamewithlimits{argmax}_{y\in Y^{*}}\left\{\sum_{j}I\left(y\in\operatornamewithlimits{argmax}\left\{s_{j}(\cdot,y^{\prime})\right\}\right)\right\}.$ (1)
In soft voting the ensemble averages the scores assigned to each response label value by the models in which case the ensemble choice aggregator will be:
$C(Y^{*},s)=\operatornamewithlimits{argmax}_{y\in Y^{*}}\left\{\frac{1}{m}\sum_{j=1}^{m}s_{j}(\cdot,y)\right\}.$ (2)
In either case $C(\cdot,s)$ is nondegenerate and satisfies unanimity, and so the ECD theorem implies a violation of model choice reversal in multiclass or regression problems. Relating this back to our random forest motivating example, the use of a unit vote for the most popular class [e.g. Breiman, 2001] implemented as a form of soft voting implies a violation of model choice reversal; the decision making inconsistency that we used to motivate our argument is indeed ubiquitous.
In stacked generalization we train a collection of models on a data set and then learn an aggregation method that assembles the output of these models into a single prediction as a secondary learning task [Wolpert, 1992]. One approach to this would be to learn a set of weights that could be used to aggregate the individual models. For example, in the case where we have a set of Bayesian regression models combined via learned weights the posterior distribution over the parameters of each model induces a score for each $y\in Y$ when given $x$ in the form of the model evidence $p_{j}(y|x)$. Thus the model scoring function is $s_{j}(x,y)=p_{j}(y|x)$ and the ensemble choice aggregator is:
$C(Y^{*},s)=\operatornamewithlimits{argmax}_{y\in Y^{*}}\left\{\sum_{j}w_{j}s_{j}(x,y)\right\}$
where $w_{j}$ is the learned weight for model $j$. If $w_{j}<0$ for some $j$ then the ensemble choice aggregator will violate ensemble unanimity. If $w_{j}=0$ for all but one model then this is not a nondegenerate ensemble choice aggregator, for example as a consequence of using a model specification method, e.g. the adaptive lasso [see Zou, 2006]. If $w_{j}>0$ for all $j$ then it is a nondegenerate ensemble choice aggregator and satisfies ensemble unanimity, and so by the ECD theorem is guaranteed to violate model choice reversal. Learning simply imposes a particular violation as an accidental by-product of optimizing the loss function.
Finally, reinforcement learning methods have been introduced in which multiple agents work together in smart collaboration to maximize rewards via voting [e.g. Xu et al., 2019]. As a simple example, in $Q$-learning, agents learn to encode the long-term value for taking actions a function mapping from an action $a$ and state of the world $\sigma$, the so-called $Q$-value, $Q(\sigma,a)$. If each agent learns its own $Q$-value, these values are aggregated via soft voting, and the action associated with the largest is chosen then the ECD theorem implies a violation of model choice reversal; similar states of the world with very similar state-action values will sometimes lead to radically different decisions.
We summarize these observations in the following corollary of the ECD theorem:
Corollary 4.1.
If $C(\cdot,s)$ satisfies equations 1 or 2 then it violates model choice reversal.
Discussion
Specifying a choice aggregator as a nondegenerate ensemble and satisfying ensemble unanimity are otherwise relatively undemanding requirements and it is easy to write down score aggregators that will use the input of more than one model and respect agreement among all models; indeed a simple weighted average where all weights are positive works. Model choice reversal–the requirement that the ensemble cannot change its mind unless some of the underlying models do so–is more difficult to guarantee, and so we focus discussion on this property.
Intuitively, model choice reversal is a form of smoothnesss. When model choice reversal holds “small” changes in individual model scores cannot provoke “large” changes in aggregated decisions. On the other hand when model choice reversal is violated the implication is that the ensemble cannot guarantee that 1) small changes to model scoring functions via training updates or 2) two nearly identical inputs will generate similar decisions. Small changes to the scores will generate large changes to the outputs in at least some part of the scoring space; thus aggregation of multiple models must necessarily introduce “discontinuities” in the decisions made by an ensemble. Inability to guarantee equal treatment of nearly equivalent cases seems to violate basic fairness requirements. We suspect that this will manifest as unpredictable and unintuitive behavior that complicates the explainability of an ensemble and compromises the user experience for a non-expert.
For example, model choice reversal is closely related to the idea of monotonicity:
Definition 4.2.
Say that $C(\cdot,s)$ is monotonic if for any $y,y^{\prime}\in Y$ and any $s,s^{\prime}\in S(Y)$ such that $s_{j}(\cdot,y)\geq s_{j}(\cdot,y^{\prime})$ implies $s^{\prime}_{j}(\cdot,y)\geq s_{j}^{\prime}(\cdot,y^{\prime})$ for all $j$ we have that $y\in C(\{y,y^{\prime}\},s)$ implies that $y\in C(\{y,y^{\prime}\},s^{\prime})$.
A monotonic aggregator is one in which more individual model support for a particular label will not make the ensemble less likely to choose it. It is straightforward to show that any monotonic ensemble score aggregator will respect model choice reversal (see Lemma 5.2 in the appendix). The following corollary is then an immediate consequence of our ECD theorem:
Corollary 4.3.
When $|Y|\geq 3$ then if $C(\cdot,s)$ is a nondegenerate ensemble choice aggregator that satisfies insertion/deletion consistency and ensemble unanimity it violates monotonicity.
So, where model choice reversal is violated monotonicity will also be violated and there will be sets of scores–and inputs associated with them–where more individual model support for a label will rule that label out as an ensemble choice, an intrinsic property of an ensemble that downstream users may find difficult to grasp and so one that undermines their trust.
The inconsistency of model choice reversal with the other properties is also tied to the capacity of the individual models within the ensemble to represent the relationship between features and labels. We observe that in our setting we can think about formalizing model representational capacity in terms of two properties. First, the range of the scoring function, $S_{j}(Y)$, bounds the ability of the model to encode information about the label. When we defined this range above we required that it include scores consistent with any ordering of the label values. Second, if each of these scores is mapped to from some point in the feature space–if $s_{j}(\cdot)$ is surjective–then we can conclude that the model has enough capacity to represent any relationship between features and labels.
This situation is sufficient to guarantee that violations of the properties in our ECD theorem are “observable” in the sense that they can actually be encountered at inference because real points in feature space lead to the scores associated with violations of definitions 2.2, 2.3, and 2.4. Restating this discussion formally:
Corollary 4.4.
Suppose that $|Y|\geq 3$ and that $C(\cdot,s)$ is a nondegenerate ensemble choice aggregator. If $s(\cdot)$ is surjective then there exists $X^{*}\subseteq X$ such that for all $x^{*}\in X^{*}$ we have that $C(\cdot,s(x^{*}))$ violates at least one of insertion/deletion consistency, ensemble unanimity, or model choice reversal.
High capacity individual models increase the risk of violations of at least one of the properties analyzed here and–to the extent that the other properties are easily satisfied–to violations of model choice reversal. It follows that there is a fundamental tradeoff between the flexibility of the individual models in an ensemble and violations of model choice reversal.
This also suggests that restrictions to the ranges of individual models or to the mapping from feature space to scores may eliminate portions of the scoring space that violate model choice reversal. For example, in the weighted averaging of Bayesian models discussed above if the distributional assumptions of the individual Bayesian regressions impose a unimodal structure on the model evidence this would amount to a restriction on the range of the individual models and so on the domain of the aggregator; under this restricted domain using majority rule as an ensemble choice aggregator would satisfy our three properties [see Black, 1948]. It also strikes us that the requirements on the model scoring function range and the surjectivity of $s_{j}(\cdot)$ for Corollary 4.4 are likely much stronger assumptions in regression problems than in classification problems; though the ECD theorem holds in a technical sense in regression we speculate that the risk of running into model choice reversal violations is lower in regression problems.
5 Conclusion
In this note we have analyzed a set of four desirable properties on ensemble information use and prediction stability and shown that at least one must be violated in any machine learning ensemble, establishing a fundamental tradeoff between information use and prediction instability. Any ensemble that uses the insights of multiple models will make choices in a counterintuitive or unstable fashion at least some of the time. Since three of these four properties are fairly easy to satisfy the last–model choice reversal–must fail in many common use cases thus introducing ‘discontinuities’ into the predictions and decisions made by ensembles. Small changes in underlying model scores can lead to large, substantive changes in ensemble predictions and choices; the ensemble can change its prediction or choice even if none of the underlying models do and there is no guarantee that similar feature space inputs will be treated similarly.
We observe that the ECD theorem presented here shows that all ensembles will behave badly some of the time with respect to the prediction stability properties we have analyzed. It seems intuitive that the risk of prediction or choice instability depends on the number of models in the underlying ensemble and on the relationship between the size of the data sample used to train the ensemble and the population it is drawn from. Accordingly we aim to extend the analysis here to investigate asymptotic ensemble behavior when the number of models in the ensemble grows large. We conjecture that the Borda count method of aggregation may satisfy definitions 2.1, 2.2, and 2.3 with a slight weakening of model choice reversal in the asymptotic case of many models. We also conjecture that as the data available for training grows arbitrarily large this will act as an implicit restriction on the diversity of the models in the ensemble and so minimize the risk of prediction instability. Finally, we conjecture that the argument developed here would apply to using a data sampling process like bootstrapping or $k$-fold cross validation to choose among multiple model architectures during training.
Acknowledgements
This work was supported by the exceptional generosity of the PPC Foundation, Inc. Additionally, we would like to thank Ernesto Guerra–Vallejos, RJ Nowling, David Andrzejewski, and the audience of the Medical College of Wisconsin SEAWINDS Data Science Symposium for helpful comments, suggestions, and feedback. Finally, we thank Christopher Taylor and Sheila Ross for their support.
Appendix
Lemma 5.1.
If $C(\cdot,s)$ satisfies transitivity then:
  1. 1.
    $y\in C(\{y,y^{\prime}\},s)$ and $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$ implies that $y=C(\{y,y^{\prime\prime}\},s)$;
  2. 2.
    $y=C(\{y,y^{\prime}\},s)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$ implies that $y=C(\{y,y^{\prime\prime}\},s)$;
  3. 3.
    $y=C(\{y,y^{\prime}\},s)$ and $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$ implies that $y=C(\{y,y^{\prime\prime}\},s)$.
Proof.
Suppose that $C(\cdot)$ satisfies transitivity. To see part 1) assume $y\in C(\{y,y^{\prime}\},s)$ and $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$. Then by transitivity $y\in C(\{y,y^{\prime\prime}\},s)$. Suppose that $y^{\prime\prime}\in C(\{y,y^{\prime\prime}\},s)$. But then by transitivity $y^{\prime\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$, a contradiction. To see part 2) assume $y=C(\{y,y^{\prime}\},s)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$. Then by transitivity $y\in C(\{y,y^{\prime\prime}\},s)$. Suppose that $y^{\prime\prime}\in C(\{y,y^{\prime\prime}\},s)$. But then by transitivity $y^{\prime}\in C(\{y,y^{\prime}\},s)$, a contradiction. To see, part 3) assume that $y=C(\{y,y^{\prime}\},s)$ and $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$ but that $y^{\prime\prime}\in C(\{y,y^{\prime\prime}\},s)$. Then by part 1) we have $y^{\prime\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$, a contradiction. ∎
Proof of Lemma 3.2.
Proof.
Sufficiency. Suppose that $C(\cdot)$ is transitive. Take any $s\in S(Y)$, any $Y_{1}^{*},Y_{2}^{*}\subseteq Y$ such that $Y_{1}^{*}\subseteq Y_{2}^{*}$. Suppose that $Y_{1}^{*}\bigcap C(Y_{2}^{*},s)\neq\emptyset$. Take any $y\in C(Y_{1}^{*},s)$; we wish to show that $y\in Y_{1}^{*}\bigcap C(Y_{2}^{*},s)$. Clearly $y\in Y_{1}^{*}$. Suppose that $y\notin C(Y_{2}^{*},s)$. Take any $y^{\prime}\in Y_{1}^{*}\bigcap C(Y_{2}^{*},s)$. Then $y^{\prime}\in C(Y_{2}^{*},s)$ implies that $y^{\prime}\in C(,s)$, and $y\in C(Y_{1}^{*},s)$ implies that $\{y,y^{\prime}\}=C(\{y,y^{\prime}\},s)$. Since $y\notin C(Y_{2}^{*},s)$ it follows that there exists $y^{\prime\prime}\in Y_{2}^{*}$ such that $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. But then by transitivity and Lemma 5.1 it must be that $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$, a contradiction. Next take any $y\in Y_{1}^{*}\bigcap C(Y_{2}^{*},s)$. Then there does not exist $y^{\prime}\in Y_{2}^{*}$ such that $y^{\prime}=C(\{y,y^{\prime}\},s)$. Thus there does not exist $y^{\prime}\in Y_{1}^{*}$ such that $y^{\prime}=C(\{y,y^{\prime}\},s)$ and so $y\in C(Y_{1}^{*},s)$. It follows that $C(Y_{1}^{*},s)=Y_{1}^{*}\bigcap C(Y_{2}^{*},s)$ and so definition 2.1 holds.
Necessity. Suppose that $C(\cdot)$ satisfies definition 2.1 but that transitivity does not hold. Then there exists $y,y^{\prime},y^{\prime\prime}$ such that $y\in C(\{y,y^{\prime}\},s)$ and $y^{\prime}\in C(\{y^{\prime},y^{\prime\prime}\},s)$, yet $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. There are four cases to consider.
Case 1: $\{y,y^{\prime}\}=C(\{y,y^{\prime}\},s)$, $\{y^{\prime},y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},s)$, and $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. By definition of $C(\cdot)$ we have $C(\{y,y^{\prime},y^{\prime\prime}\},s)=\{y^{\prime},y^{\prime\prime}\}$. But then we have $\{y,y^{\prime}\}\subseteq\{y,y^{\prime},y^{\prime\prime}\}$ and $\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)\neq\emptyset$ and yet $C(\{y,y^{\prime}\},s)=\{y,y^{\prime}\}\neq y^{\prime}=\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)$, a violation of definition 2.1 and so a contradiction.
Case 2: $\{y,y^{\prime}\}=C(\{y,y^{\prime}\},s)$, $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$, and $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. By definition of $C(\cdot)$ we have $C(\{y,y^{\prime},y^{\prime\prime}\},s)=y^{\prime}$. But then we have $\{y,y^{\prime}\}\subseteq\{y,y^{\prime},y^{\prime\prime}\}$ and $\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)\neq\emptyset$ and yet $C(\{y,y^{\prime}\},s)=\{y,y^{\prime}\}\neq y^{\prime}=\{y,y^{\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)$, a violation of definition 2.1 and so a contradiction.
Case 3: $y=C(\{y,y^{\prime}\},s)$, $\{y^{\prime},y^{\prime\prime}\}=C(\{y^{\prime},y^{\prime\prime}\},s)$, and $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. By definition of $C(\cdot)$ we have $C(\{y,y^{\prime},y^{\prime\prime}\},s)=y^{\prime\prime}$. But then we have $\{y^{\prime},y^{\prime\prime}\}\subseteq\{y,y^{\prime},y^{\prime\prime}\}$ and $\{y^{\prime},y^{\prime\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)\neq\emptyset$ and yet $C(\{y^{\prime},y^{\prime\prime}\},s)=\{y^{\prime},y^{\prime\prime}\}\neq y^{\prime\prime}=\{y^{\prime},y^{\prime\prime}\}\bigcap C(\{y,y^{\prime},y^{\prime\prime}\},s)$, a violation of definition 2.1 and so a contradiction.
Case 4: $y=C(\{y,y^{\prime}\},s)$, $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$, and $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. But then $C(\{y,y^{\prime},y^{\prime\prime}\},s)=\emptyset$, a contradiction. ∎
Proof of Lemma 3.3.
Proof.
It suffices to consider two cases. First suppose that $C(\{y,y^{\prime}\},s)=\{y,y^{\prime}\}$ but that without loss of generality $C(\{y,y^{\prime}\},s^{\prime})=\{y\}$. Then there is an immediate contradiction of model choice reversal, since there does not exist $j$ such that $s_{j}(\cdot,y^{\prime})>s_{j}(\cdot,y)$ and $s_{j}^{\prime}(\cdot,y^{\prime})<s_{j}^{\prime}(\cdot,y)$ and so the claim holds. Next suppose without loss of generality that $C(\{y,y^{\prime}\},s)=\{y\}$ but that $y^{\prime}\in C(\{y,y^{\prime}\},s^{\prime})$. But then again there is an immediate contradiction of model choice reversal, since there does not exist $j$ such that $s_{j}(\cdot,y^{\prime})<s_{j}(\cdot,y)$ and $s_{j}^{\prime}(\cdot,y^{\prime})>s_{j}^{\prime}(\cdot,y)$ and so the claim holds. Thus $C(\{y,y^{\prime}\},s)=C(\{y,y^{\prime}\},s^{\prime})$. ∎
Proof of Theorem Theorem.
Proof.
Suppose that $|Y|\geq 3$, that $C(\cdot,s)$ satisfies definition 2.2, and also by contradiction that it satisfies all of definitions 2.1, 2.3, and 2.4. We will prove the result in a series of steps, following the basic structure of Fey (2014). To begin, note that by Lemma 3.2 $C(\cdot,s)$ satisfies definition 3.1. Say that $j$ is decisive for $y$ over $y^{\prime}$ if $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ implies that $y\in C(\{y,y^{\prime}\},s)$.Take any $y,y^{\prime}\in Y$, $y\neq y^{\prime}$.
Step 1: identify a critical model $\mathbf{j^{*}}$. Consider the following two ensemble scores $s,s^{\prime}\in S(Y)$:
$\begin{array}[]{llll}s_{1}(\cdot,y)&>s_{1}(\cdot,y^{\prime})&>\ldots\\ \vdots&\vdots&\ddots\\ s_{m}(\cdot,y)&>s_{m}(\cdot,y^{\prime})&>\ldots\end{array}$
and:
$\begin{array}[]{llll}s^{\prime}_{1}(\cdot,y^{\prime})&>s^{\prime}_{1}(\cdot,y)&>\ldots\\ \vdots&\vdots&\ddots\\ s^{\prime}_{m}(\cdot,y^{\prime})&>s^{\prime}_{m}(\cdot,y)&>\ldots\end{array}$
By ensemble unanimity, $y=C(\{y,y^{\prime}\},s)$ and $y^{\prime}=C(\{y,y^{\prime}\},s^{\prime})$. Now, consider a set of scores $s^{j}$ such that:
$\begin{array}[]{llll}s^{j}_{1}(\cdot,y^{\prime})&>s^{j}_{1}(\cdot,y)&>\ldots\\ \vdots&\vdots&\ddots\\ s^{j}_{j}(\cdot,y^{\prime})&>s^{j}_{j}(\cdot,y)&>\ldots\\ s^{j}_{j+1}(\cdot,y)&>s^{j}_{j+1}(\cdot,y^{\prime})&>\ldots\\ \vdots&\vdots&\ddots\\ s^{j}_{m}(\cdot,y)&>s^{j}_{m}(\cdot,y^{\prime})&>\ldots\end{array}$
Define $j^{*}$ such that $y\in C(\{y,y^{\prime}\},s^{j^{*}-1})$ and $y^{\prime}\in C(\{y,y^{\prime}\},s^{j^{*}})$.
Step 2: for all $\mathbf{y^{\prime\prime}\neq y,y^{\prime}}$ we have $\mathbf{j^{*}}$ decisive for $\mathbf{y^{\prime}}$ over $\mathbf{y^{\prime\prime}}$. To see this, take any $y^{\prime\prime}$ and consider $s$ such that:
$\begin{array}[]{lllll}s_{1}(\cdot,y^{\prime})&>s_{1}(\cdot,y^{\prime\prime})&>s_{1}(\cdot,y)&\ldots\\ \vdots&\vdots&\vdots&\ddots\\ s_{j^{*}-1}(\cdot,y^{\prime})&>s_{j^{*}-1}(\cdot,y^{\prime\prime})&>s_{j^{*}-1}(\cdot,y)&\ldots\\ s_{j^{*}}(\cdot,y)&>s_{j^{*}}(\cdot,y^{\prime})&>s_{j^{*}}(\cdot,y^{\prime\prime})&\ldots\\ \vdots&\vdots&\vdots&\ddots\\ s_{m}(\cdot,y)&>s_{m}(\cdot,y^{\prime})&>s_{m}(\cdot,y^{\prime\prime})&\ldots\\ \end{array}$ (9)
Note that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ if and only if $s^{j^{*}-1}_{j}(\cdot,y)>s^{j^{*}-1}_{j}(\cdot,y^{\prime})$ for all $j$ and so by Lemma 3.3 $y\in C(\{y,y^{\prime}\},s)$. By ensemble unanimity $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$. Thus by Lemma 5.1 $y=C(\{y,y^{\prime\prime}\},s)$. Next consider the set of scores $S^{*}\subseteq S(Y)$ such that for all $s^{*}\in S^{*}$:
$\begin{array}[]{lll}s^{*}_{1}(y^{\prime})\backslash s^{*}_{1}(y^{\prime\prime})&>s^{*}_{1}(y)&>\ldots\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,y^{\prime})&>s^{*}_{j^{*}}(\cdot,y)&>s^{*}_{j^{*}}(\cdot,y^{\prime\prime})\ldots\\ s^{*}_{j^{*}+1}(\cdot,y)&>s^{*}_{j^{*}+1}(\cdot,y^{\prime})&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{j^{*}+1}(\cdot,y^{\prime\prime})&\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,y)&>s^{*}_{m}(\cdot,y^{\prime})&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{m}(\cdot,y^{\prime\prime})&\end{array}$ (18)
where $s(a)\backslash s(b)$ denotes any of $s(\cdot,a)>s(\cdot,b)$, $s(\cdot,a)<s(\cdot,b)$ or $s(\cdot,a)=s(\cdot,b)$. Note that for any $s^{*}\in S^{*}$ we have $s^{*}_{j}(\cdot,y)>s_{j}^{*}(\cdot,y^{\prime})$ if and only if $s_{j}^{j^{*}}(\cdot,y)>s_{j}^{j^{*}}(\cdot,y^{\prime})$ for all $j$ and so by Lemma 3.3 we have $y^{\prime}\in C(\{y,y^{\prime}\},s^{*})$. Moreover, for any $s^{*}\in S^{*}$ we have that $s^{*}_{j}(\cdot,y)>s_{j}^{*}(\cdot,y^{\prime\prime})$ if and only if the same holds in 9 for all $j$ and so Lemma 3.3 also implies that $y=C(\{y,y^{\prime\prime}\},s^{*})$. Thus by Lemma 5.1 we have $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,y^{\prime})>s_{j^{*}}(\cdot,y^{\prime\prime})$ there exists $s^{*}\in S^{*}$ such that $s_{j}(\cdot,y^{\prime})>s_{j}(\cdot,y^{\prime\prime})$ if and only if $s^{*}_{j}(\cdot,y^{\prime})>s^{*}_{j}(\cdot,y^{\prime\prime})$. By Lemma 3.3 we have $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$. We conclude that $j^{*}$ is decisive for $y^{\prime}$ over $y^{\prime\prime}$.
Step 3: for all $\mathbf{y^{\prime\prime}\neq y,y^{\prime}}$ we have $\mathbf{j^{*}}$ decisive for $\mathbf{y}$ over $\mathbf{y^{\prime\prime}}$. Consider the set of ensemble scores $S^{*}\subseteq S(Y)$ such that for all $s^{*}\in S^{*}$:
$\begin{array}[]{lll}s^{*}_{1}(\cdot,y)\backslash s^{*}_{1}(\cdot,y^{\prime\prime})&>s^{*}_{1}(\cdot,y^{\prime})&>\ldots\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,y)&>s^{*}_{j^{*}}(\cdot,y^{\prime})&>s^{*}_{j^{*}}(\cdot,y^{\prime\prime})>\ldots\\ s^{*}_{j^{*}+1}(\cdot,y)&>s^{*}_{j^{*}+1}(\cdot,y^{\prime})&>\ldots\\ \hskip 19.91692pt\backslash s^{*}_{j^{*}+1}(\cdot,y^{\prime\prime})&&\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,y)\backslash s^{*}_{m}(\cdot,y^{\prime\prime})&>s^{*}_{m}(\cdot,y^{\prime})&>\ldots\\ \end{array}$ (26)
By step (2) we have that $j^{*}$ is decisive for $y^{\prime}$ over $y^{\prime\prime}$ and so $y^{\prime}=C(\{y^{\prime},y^{\prime\prime}\},s^{*})$. By ensemble unanimity we have $y=C(\{y,y^{\prime}\},s^{*})$. By Lemma 5.1 we have $y=C(\{y,y^{\prime\prime}\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,y)>s_{j^{*}}(\cdot,y^{\prime\prime})$ there exists $s^{*}\in S^{*}$ such that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime\prime})$ if and only if $s^{*}_{j}(\cdot,y)>s_{j}^{*}(\cdot,y^{\prime\prime})$ for all $j$ and so by Lemma 3.3 we have that $y=C(\{y,y^{\prime\prime}\},s)$. We conclude that $j^{*}$ is decisive for $y$ over $y^{\prime\prime}$.
Step 4: for all $\mathbf{y^{\prime\prime}\neq y,y^{\prime}}$ we have $\mathbf{j^{*}}$ decisive for $\mathbf{y^{\prime\prime}}$ over $\mathbf{y}$. Consider $s$ such that:
$\begin{array}[]{lllll}s_{1}(\cdot,y^{\prime})&>s_{1}(\cdot,y^{\prime\prime})&>s_{1}(\cdot,y)&\ldots\\ \vdots&\vdots&\vdots&\ddots\\ s_{j^{*}}(\cdot,y^{\prime\prime})&>s_{j^{*}}(\cdot,y)&>s_{j^{*}}(\cdot,y^{\prime})&\ldots\\ \vdots&\vdots&\vdots&\ddots\\ s_{m}(\cdot,y^{\prime\prime})&>s_{m}(\cdot,y)&>s_{m}(\cdot,y^{\prime})&\ldots\\ \end{array}$ (32)
Note that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ iff $s^{j^{*}-1}_{j}(\cdot,y)>s^{j^{*}-1}_{j}(\cdot,y^{\prime})$ for all $j$ and so by Lemma 3.3 we have $y\in C(\{y,y^{\prime}\},s)$. By ensemble unanimity $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$ and so by Lemma 5.1 we have $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$. Consider the set of ensemble scores $S^{*}\subseteq S(Y)$ such that for all $s^{*}\in S^{*}$:
$\begin{array}[]{lll}s^{*}_{1}(\cdot,y^{\prime})&>s^{*}_{1}(\cdot,y)&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{1}(\cdot,y^{\prime\prime})&\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,y^{\prime\prime})&>s^{*}_{j^{*}}(\cdot,y^{\prime})&>s^{*}_{j^{*}}(\cdot,y)\geq\ldots\\ s^{*}_{j^{*}+1}(\cdot,y)&>s^{*}_{j^{*}+1}(\cdot,y^{\prime})&>\ldots\\ \hskip 19.91692pt\backslash s^{*}_{j^{*}+1}(\cdot,y^{\prime\prime})&&\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,y)&>s^{*}_{m}(\cdot,y^{\prime})&>\ldots\\ \hskip 19.91692pt\backslash s^{*}_{m}(\cdot,y^{\prime\prime})&&\end{array}$ (42)
Note that for any $s^{*}\in S^{*}$ we have $s^{*}_{j}(\cdot,y)>s^{*}_{j}(\cdot,y^{\prime})$ iff $s^{j^{*}}_{j}(\cdot,y)>s^{j^{*}}_{j}(\cdot,y^{\prime})$ and so by Lemma 3.3 we have $y^{\prime}\in C(\{y,y^{\prime}\},s^{*})$. Also, $s^{*}_{j}(\cdot,y^{\prime})>s^{*}_{j}(\cdot,y^{\prime\prime})$ iff the same holds in 32 and so by Lemma 3.3 we have $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s^{*})$. By Lemma 5.1 it follows that $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,y^{\prime\prime})>s_{j^{*}}(\cdot,y)$ there exists $s^{*}\in S^{*}$ such that $s_{j}(\cdot,y^{\prime\prime})>s_{j}(\cdot,y)$ iff $s_{j}^{*}(\cdot,y^{\prime\prime})>s_{j}^{*}(\cdot,y)$ for all $j$ and so by Lemma 3.3 we have $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. We conclude that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y$.
Step 5: for all $\mathbf{y^{\prime\prime}\neq y,y^{\prime}}$ we have $\mathbf{j^{*}}$ decisive for $\mathbf{y^{\prime\prime}}$ over $\mathbf{y^{\prime}}$. Consider the set of ensemble scores $S^{*}\subseteq S(Y)$ such that for all $s^{*}\in S^{*}$:
$\begin{array}[]{lll}s^{*}_{1}(\cdot,y)&>s^{*}_{1}(\cdot,y^{\prime})&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{1}(\cdot,y^{\prime\prime})&\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,y^{\prime\prime})&>s^{*}_{j^{*}}(\cdot,y)&>s^{*}_{j^{*}}(\cdot,y^{\prime})>\ldots\\ s^{*}_{j^{*}+1}(\cdot,y)&>s^{*}_{j^{*}+1}(\cdot,y^{\prime})&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{j^{*}+1}(\cdot,y^{\prime\prime})&\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,y)&>s^{*}_{m}(\cdot,y^{\prime})&>\ldots\\ &\hskip 19.91692pt\backslash s^{*}_{m}(\cdot,y^{\prime\prime})&\end{array}$ (52)
By step (4) we have that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y$, and so $y^{\prime\prime}=C(\{y,y^{\prime\prime}\},s)$. By ensemble unanimity $y=C(\{y,y^{\prime}\},s^{*})$. By Lemma 5.1 we have $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,y^{\prime\prime})>s_{j^{*}}(\cdot,y^{\prime})$ there exists $s^{*}\in S^{*}$ such that $s_{j}(\cdot,y^{\prime\prime})>s_{j}(\cdot,y^{\prime})$ iff $s^{*}_{j}(\cdot,y^{\prime\prime})>s^{*}_{j}(\cdot,y^{\prime})$ for all $j$ and so by Lemma 3.3 we have that $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s)$. We conclude that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y^{\prime}$.
Step 6: $\mathbf{j^{*}}$ decisive for $\mathbf{y}$ over $\mathbf{y^{\prime}}$ and for $\mathbf{y^{\prime}}$ over $\mathbf{y}$. Consider the set of ensemble scores $S^{*}\subseteq S(Y)$ such that for all $s^{*}\in S^{*}$:
$\begin{array}[]{lll}s^{*}_{1}(\cdot,y)&>&\ldots\\ \hskip 19.91692pt\backslash s^{*}_{1}(\cdot,y^{\prime})&&\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,y)&>s^{*}_{j^{*}}(\cdot,y^{\prime\prime})&>s^{*}_{j^{*}}(\cdot,y^{\prime})>\ldots\\ s^{*}_{j^{*}+1}(\cdot,y)&>&\ldots\\ \hskip 19.91692pt\backslash s^{*}_{j^{*}+1}(\cdot,y^{\prime})&&\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,y)&>&\ldots\\ \hskip 19.91692pt\backslash s^{*}_{m}(\cdot,y^{\prime})&&\end{array}$ (62)
By step (5) we have that $j^{*}$ is decisive for $y^{\prime\prime}$ over $y^{\prime}$ and so $y^{\prime\prime}=C(\{y^{\prime},y^{\prime\prime}\},s^{*})$. By step (3) we have that $j^{*}$ is decisive for $y$ over $y^{\prime\prime}$ and so $y=C(\{y,y^{\prime\prime}\},s^{*})$. By Lemma 5.1 we have $y=C(\{y,y^{\prime}\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,y)>s_{j^{*}}(\cdot,y^{\prime})$ there exists $s^{\prime}\in S^{*}$ such that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ iff $s^{*}_{j}(\cdot,y)>s^{*}_{j}(\cdot,y^{\prime})$ and so by Lemma 3.3 we have $y=C(\{y,y^{\prime}\},s)$. We conclude that $j^{*}$ is decisive for $y$ over $y^{\prime}$. The argument for $j^{*}$ decisive for $y^{\prime}$ over $y$ is the same with the two reversed, citing step (4) and step (2).
Step 7: $\mathbf{j^{*}}$ is decisive for any $a\in Y$ over any $b\in Y$. Take any $a\neq y,y^{\prime}$ and any $b\neq y,y^{\prime}$ and consider the set of scores $S^{*}\subseteq S(Y)$ such that for all $s^{*}\in S^{*}$:
$\begin{array}[]{lll}s^{*}_{1}(\cdot,a)&>s^{*}_{1}(\cdot,y)&\ldots\\ \hskip 19.91692pt\backslash s^{*}_{1}(\cdot,b)&\hskip 19.91692pt\backslash s^{*}_{1}(\cdot,y^{\prime})&\\ \vdots&\vdots&\ddots\\ s^{*}_{j^{*}}(\cdot,a)&>s^{*}_{j^{*}}(\cdot,y)&>s^{*}_{j^{*}}(\cdot,y^{\prime})>s^{*}_{j^{*}}(\cdot,b)\ldots\\ \vdots&\vdots&\ddots\\ s^{*}_{m}(\cdot,a)&>s^{*}_{m}(\cdot,y)&\ldots\\ \hskip 19.91692pt\backslash s^{*}_{m}(\cdot,b)&\hskip 19.91692pt\backslash s^{*}_{m}(\cdot,y^{\prime})&\end{array}$ (70)
By step (4) we have that $j^{*}$ is decisive for $a$ over $y$, by step (6) we have that $j^{*}$ decisive for $y$ over $y^{\prime}$, and by step (2) we have that $j$ is decisive for $y^{\prime}$ over $b$. Thus by Lemma 5.1 we have $a=C(\{a,b\},s^{*})$. But now note that for any $s$ such that $s_{j^{*}}(\cdot,a)>s_{j^{*}}(\cdot,b)$ there exists $s^{*}\in S^{*}$ such that $s_{j}(\cdot,a)>s_{j}(\cdot,b)$ iff $s^{*}_{j}(\cdot,a)>s^{*}_{j}(\cdot,b)$ for all $j$, and so by Lemma 3.3 we have $a=C(\{a,b\},s)$. Thus $j^{*}$ decisive for $a$ over $b$ and since $a$ and $b$ were arbitrary, there exists $j$, namely $j^{*}$, such that for all $y,y^{\prime}\in Y$ if $s_{j^{*}}(\cdot,y)>s_{j^{*}}(\cdot,y^{\prime})$ then $y=C(\{y,y^{\prime}\},s)$. Thus definition 2.2 is violated, a contradiction. ∎
Lemma 5.2.
If $C(\cdot,s)$ is monotonic then it respects model choice reversal.
Proof.
Take any $C(\cdot)$ and suppose that it is monotonic. Take any $y,y^{\prime}\in Y$ and any $s,s^{\prime}$ such that $y\in C(\{y,y^{\prime}\},s)$, $y^{\prime}\notin C(\{y,y^{\prime}\},s)$, and $y^{\prime}\in C(\{y,y^{\prime}\},s^{\prime})$. Suppose by contradiction that there does not exist $j$ such that $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ and $s^{\prime}_{j}(\cdot,y)<s_{j}^{\prime}(\cdot,y^{\prime})$. That is, for all $j$ if $s_{j}(\cdot,y)>s_{j}(\cdot,y^{\prime})$ then $s_{j}^{\prime}(\cdot,y)\geq s_{j}^{\prime}(\cdot,y^{\prime})$ and if $s_{j}^{\prime}(\cdot,y)<s_{j}^{\prime}(\cdot,y^{\prime})$ then $s_{j}(\cdot,y)\geq s_{j}(\cdot,y^{\prime})$. But then $s_{j}^{\prime}(\cdot,y^{\prime})\geq s_{j}^{\prime}(\cdot,y)$ implies that $s_{j}(\cdot,y^{\prime})\geq s_{j}(\cdot,y)$ for all $j$. By monotonicity $y^{\prime}\in C(\{y,y^{\prime}\},s^{\prime})$ implies that $y^{\prime}\in C(\{y,y^{\prime}\},s)$, a contradiction. ∎
Proof of Corollary 4.4.
Proof.
Suppose that $|Y|\geq 3$, that $C(Y^{*},s)\neq\emptyset$ for all nonempty $Y^{*}\subseteq Y$, and that $C(\cdot,s)$ is insertion/deletion consistent. Then by the ECD theorem there exists $S^{*}\subseteq S(Y)$ such that at least one of definitions 2.2, 2.3, or 2.4 are violated on $S^{*}$. By $s(\cdot)$ surjective, for all $s^{*}\in S^{*}$ there exists $x\in X$ such that $s(x)=s^{*}$. Thus, define:
$X^{*}\equiv\bigcup_{s^{*}\in S^{*}}\left\{x\in X|s(x)=s^{*}\right\}.$